OpenStudy (vortish):

3/8=1/2 x do you find the common denominator and then cross multiply for this help

OpenStudy (anonymous):

You are trying to find x, correct?

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):

Step 1. Multiply both sides by 2

hero (hero):

3/8 = x/2 Cross multiply 6 = 8x 6/8 = x 3/4 = x

OpenStudy (anonymous):

If you equation is (3/8)=(1/2)x then multiply both sides by 2 and you'll get your answer for x

OpenStudy (vortish):

ok i come up with x = 6/8 and 3/4 so if we have two fractions which is correct or are they both correct

OpenStudy (anonymous):

They would be both correct. They are both identical